# Emergent morphogenesis via planar fabrication enabled by a reduced model of composites

### v1

@@ II.2.2 Energy Formulation for Stimulus-Responsive Composite Shells
$\begin{gathered}\varepsilon_{i}^{\mathrm{th}}=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{d_{i}}{d_{\max}},\quad d_{i}=\min_{j\in\mathcal{B}}\bigl{\lVert}\mathbf{x}_{i}-\mathbf{x}_{j}\bigr{\rVert},\quad d_{\max}=\max_{i\in\Omega\setminus\mathcal{B}}d_{i},\end{gathered}$ (7)
@@
where $\Omega$ is the set of all edge midpoints in the composite domain. $\mathcal{B}$ is the subset of $\Omega$ lying inside the PLA and Shrinky Dink bilayer region. $\mathbf{x}_{i}$ is the midpoint coordinate of edge $i\in\Omega\setminus\mathcal{B}$, and $\mathbf{x}_{j}$ of edge $j\in\mathcal{B}$. $\varepsilon_{pre}^{\mathrm{th}}$ is the thermal strain that develops in a pure Shrinky Dink sheet. $d_{i}$ is the shortest distance from midpoint $i$ to the PLA and Shrinky Dink bilayer region. $d_{\max}$ is the largest distances $d_{i}$ over the region.
@@
$\Delta\varepsilon_{j}(\bm{X},T)=\sum_{p=1}^{4}s_{p}[\varepsilon_{jp}(\bm{X},T)-\varepsilon^{th}_{jp}(\bm{X},T)],$ (8)
@@
II.3 Numerical Results Validated with Experiments
The bilayer composite is modeled as a single-layer elastic shell that incorporates thermal strains and dihedral angles as reference strains via the energy, Equation 6. Each region is characterized by its Young’s modulus $Y$. At room temperature, for the Shrinky Dink sheet, denoted as layer 1, $Y_{1}$ typically ranges from 1GPa to 3GPa mishra_effect_2005 . For the 3D-printed PLA layer, denoted as layer 2, $Y_{2}$ ranges from 2GPa to 4GPa crupano_investigating_2024 . However, at a temperature over $300^{\circ}F$ which is the requirement of the thermal actuation, the modulus becomes around two to three orders of magnitude smaller than their values at room temperature lendvai2023valorization ; tabi2019application ; ricarte2024tutorial . So we set $Y_{1}=1$MPa for the Shrinky Dink, and $Y_{2}=3$MPa for the PLA layer. Given the scaling of stiffness and energy are both linear in Young’s modulus, the unit from GPa to MPa does not affect the solution to equilibrium. The resulting effective stretching stiffness is ugural2011advanced
@@ II.3 Numerical Results Validated with Experiments
$D/b=D_{\rm eff}/b=\sum_{i=1}^{2}Y_{i}\biggl{[}\frac{h_{i}^{3}}{12}+h_{i}(y_{i}-\bar{y})^{2}\biggr{]}.$ (10)
Set the bottom of Shrinky Dink (layer 1) at $y=0$, we can define the following for each layer
$y_{1}=\frac{h_{1}}{2},\quad y_{2}=h_{1}+\frac{h_{2}}{2},\quad A_{1}=bh_{1},\quad A_{2}=bh_{2},$
and the location of the bending neutral axis is ugural2011advanced
$\bar{y}=\frac{Y_{1}A_{1}y_{1}+Y_{2}A_{2}y_{2}}{Y_{1}A_{1}+Y_{2}A_{2}}.$
@@
Recall that layer 1 in Equation 10 denotes the Shrinky Dink sheet, and layer 2 denotes the PLA layer. In our experiments, we used thicknesses of $h_{1}=0.3$mm for the Shrinky Dink sheet, and $h_{2}=0.7$ $m$, $0.6$ mm, and 1.01.0mm for the PLA layer in patterns A, B, and C in Figure 2, respectively. Since stretching is observed more difficult than bending in the experiments, we scaled the Young’s modulus for stretching, and thus $k_{s}$, up by one order of magnitude.
@@
We model the deformation as a static process, solving the equilibrium equations via the Newton-Raphson method. Thermal actuation is applied in incremental load steps with an adaptive step size chosen to satisfy the convergence criterion $|\Delta\mathbb{X}|/|\mathbb{X}|<10^{-5}$. Thermal strains are prescribed according to Equation 7, see Section II.2.2 for justification. Because the Shrinky Dink layer contracts on heating, the induced thermal strains are negative in sign, $\epsilon^{th}_{pre}<0$. To account of practical oven heating, we consider a small, spatially uniform thermal-flux load during the heating process, which decays to zero by the end of the actuation.
@@
We discretize the composite domain into triangular elements, three meshes in the reference configuration are shown in Fig. 4. We have verified the finite-difference approximations of the gradient and Hessian of the total energy, and the analytical expressions for both the gradient and Hessian of the energy are given in the Appendix. The numerical implementation was further verified with analytical solutions for uniaxial tension and compression of a simple geometry domain, and the deflection of a cantilever beam under an endpoint load. Complex behaviors, such as bulging and necking also emerge from simple deformation modes zhang_iterated_2024 . Given these verifications, all subsequent simulations are obtained with the same numerical implementations.
Figure 5: Representative snapshots, starting with the pattern of Figure 2A during the thermally actuated manufacturing process. From the top to the bottom rows, the duration of the thermal stimulus increases. In each row, (i) shows the experimentally scanned shape, (ii) shows the numerical simulation, (iii) shows the distribution of the absolute value of the dihedral angle $|\theta|$, and (iv) shows the distribution of the axial strain $\epsilon$.
@@
Figure 5 presents the experimental scans, numerical simulations, and contours of dihedral angle $\theta$ and axial strains $\epsilon$ for a circular composite with a six-arm PLA pattern, Figure 2A. The term “Stimulus” refers to the monotonically increasing stimulus, which is heat here, during the process. From top to bottom, the simulations are in equilibrium with prescribed thermal strains $\epsilon^{th}_{pre}=-0.20,-0.30$ and 0.77-0.77, respectively. As noted above, these thermal strains are nonuniform and vary with each material point’s distance to the PLA pattern. Larger thermally induced shrinkage results in more significant deformations and axial strains. A greater mismatch in axial strains leads to larger dihedral angles. To accommodate these deformations, the structure morphs from a planar configuration into a 3D shape. Deformation localizes heavily between the arms of the 3D-printed PLA pattern, where the Shrinky Dink layer undergoes the largest shrinkage, indicated by A-C(iii) and A-C(iv) in Figure 5. Overall, excellent agreement is observed between the scanned experimental shapes (first column) and the numerical simulations at each stage of the thermal actuation.
@@
Figure 6 shows the results for the rectangular composite with a two-star PLA pattern in Figure 2B, subjected to thermal strains $\epsilon^{th}_{pre}=-0.08,-0.15$ and 0.30-0.30, from the top to the bottom, respectively. As the prescribed thermal axial strain increases over time, the structure develops larger axial strains $\epsilon$, strains mismatch, dihedral angles $\theta$, and overall curvature. Deformation localizes near the four corners of the Shrinky Dink layer. The slight asymmetry in the deformed shapes arises from the mesh used in the numerical simulation and minor anisotropies in the bilayer composite due to physical manufacturing defects. Nevertheless, the numerical simulations closely approximate the scanned experimental results.
@@
Similarly, Figure 7 illustrates the deformed configurations at thermally actuated strains $\epsilon^{th}_{pre}=-0.10$, 0.30-0.30 and 0.50-0.50 for the cross pattern in Figure 2C. As the prescribed thermal strain increases, deformation concentrates increasingly between pattern’s arms. Because the Shrinky Dink layer is square rather than circular, these concentration zones form four sharp “petals” instead of the rounded bowl seen in Figure 5.
@@
These three examples demonstrate excellent agreement between scanned experimental 3D shapes and the numerical simulations morphed from 2D planar composites across various domain geometries, patterns, and levels of thermal actuation. More complex structures can be achieved by assembling multiple shapes and using multiscale modeling. The substantial computational demand of multiscale modeling across multiple time and length scales can be mitigated through iterated learning of operators for out-of-distribution data zhang_iterated_2024 . Potential applications of the three examples include deployable flower petals, bowls, frame-based pillows, canoes, and tent-type structures, each achieved by selecting an appropriate reference length scale.
II.4 Structural similarity index measure

### v2

@@ 2.2.3 Thermal Strain and Temperature Change Distribution
$d_{i}=\min_{j\in\mathcal{B}}\bigl\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\bigr\rVert,\quad d_{\max}=\max_{i\in\Omega\setminus\mathcal{B}}d_{i},$ (10)
@@
where $\Omega$ denotes the set of all edge midpoints in the composite domain, and $\mathcal{B}\subset\Omega$ denotes the subset lying within the Shrinky Dink-PLA bilayer region. Here, $\mathbf{x}_{i}$ is the midpoint coordinate of edge $i\in\Omega\setminus\mathcal{B}$, and $\mathbf{x}_{j}$ that of edge $j\in\mathcal{B}$.
To prescribe a thermal strain field consistent with experimental observations, we require that it satisfies the following conditions:
(i) $\varepsilon^{\mathrm{th}}(0)=0$ at the kirigami bilayer boundary.
(ii) $\varepsilon^{\mathrm{th}}(d_{\max})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}$, corresponding to free shrinkage far from the bilayer boundary.
(iii) $\varepsilon^{\mathrm{th}}(d)$ increases monotonically with distance $d$.
A general form satisfying these requirements is an exponential profile
$\varepsilon^{\mathrm{th}}(d)=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{1-\exp\!\left(-{d}/{L_{c}}\right)}{1-\exp\!\left(-{d_{\max}}/{L_{c}}\right)},\quad d\in[0,d_{\max}],$ (11)
where $L_{c}>0$ is a characteristic decay length controlling the spatial localization of thermal strain, and $\varepsilon_{\mathrm{pre}}^{\mathrm{th}}$ is the prescribed thermal strain parameter.
We introduce a dimensionless localization parameter $\gamma\in(0,\infty)$ defined by
$L_{c}={d_{\max}}/{\gamma}.$ (12)
With this, Equation 11 can be expressed in nondimensional form as
$\varepsilon^{\mathrm{th}}(d/d_{\max})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{1-\exp(-\gamma d/d_{\max})}{1-\exp(-\gamma)}.$ (13)
As shown by the asymptotic analysis in Appendix 5, this formulation admits two limiting cases. In the limit $\gamma\to 0$, Equation 13 reduces to
$\lim_{\gamma\to 0}\varepsilon^{\mathrm{th}}(\frac{d}{d_{\max}})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\frac{d}{d_{\max}},$ (14)
which corresponds to a linear thermal strain relation. In contrast, in the limit $\gamma\to\infty$, Equation 13 reduces to
$\lim_{\gamma\to\infty}\varepsilon^{\mathrm{th}}(\frac{d}{d_{\max}})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}},\quad d>0,$ (15)
while $\varepsilon^{\mathrm{th}}(0)=0$ for all $\gamma$. This corresponds to a strongly localized thermal strain near the kirigami bilayer boundary.
Figure 4E illustrates the variation of the normalized thermal strain $\varepsilon^{\mathrm{\rm th}}/\varepsilon_{\rm pre}^{\mathrm{\rm th}}$ with normalized distance ${d}/{d_{\max}}$ for $\gamma=0.1$, 1, and $5$, corresponding to characteristic length $L_{c}=d_{\max}/\gamma$ in Equation 12. As predicted by the asymptotic analysis, the response approaches a linear profile when $\gamma\to 0$, the case $\gamma=0.1$ is shown by black circular markers. As $\gamma$ increases, the response becomes increasingly localized near the bilayer boundary (corresponding to $d/d_{\max}=0$), as shown by the blue and red markers at $\gamma=1$ and 5, respectively. No markers are presented at $d/d_{\max}=0$ because the discrete mesh edge length is always greater than zero. Figures 4F(i)-(iii) show the deformed configurations of pattern A with $\varepsilon_{\rm pre}^{\mathrm{\rm th}}=0.7$ for $\gamma=0.1$, 1, and 5, from top to bottom, respectively. Figure 4G shows the corresponding top views of the configurations shown in Figure 4F. As $\gamma$ increases, the characteristic length $L_{c}$ decreases, leading to a more localized thermal strain field and consequently, more localized deformation, as illustrated in Figure 4F(iii) and Figure 4G(iii).
Hereafter, we adopt the linear form for thermal strain field
$\varepsilon_{i}^{\mathrm{th}}=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{d_{i}}{d_{\max}},\quad$ (16)
which corresponds to the $\gamma\to 0$ limit of the exponential mode. This choice serves as a first-order approximation that captures the experimentally observed spatial trend while avoiding the introduction of an additional length scale parameter. The varialbles $d_{i}$ and $d_{\max}$ are defined in Equation 10, and $\varepsilon_{\rm pre}^{\mathrm{th}}$ is the thermal strain that develops in a pure Shrinky Dink sheet.
@@
$\Delta\varepsilon_{j}(\bm{X},T)=\sum_{p=1}^{4}s_{p}[\varepsilon_{jp}(\bm{X},T)-\varepsilon^{\rm th}_{jp}(\bm{X},T)],$ (17)
@@
2.2.4 Model Assumptions for the Adhesive Layer
We used a two-part Gorilla epoxy adhesive to bond the Shrinky Dink and PLA kirigami layers. After curing, the thickness of the adhesive layer is in the range of $(0.0,0.1]$ mm. Recall that the thickness of the Shrinky Dink substrate is $h_{1}=0.3$mm, while the thickness of the PLA kirigami layer is in the range $h_{2}\in[0.6,1.0]$ mm for the different patterns considered. As a result, the total thickness of the bilayer composite lies in the range $[1.0,1.4]$ mm, and the adhesive layer accounts for less than 1/10 of the total thickness. Given this small thickness ratio, the adhesive contribution to the overall bending and stretching stiffness of the composite is expected to be secondary. We therefore neglect the adhesive layer in the numerical model and represent the structure as an effective bilayer composite.
If the adhesive were explicitly modeled as a third layer, reported material properties for cured epoxy adhesives indicate a Young’s modulus in the range 1.5-4.0 GPa [29, 41] at room temperature, and a Poisson’s ratio between 0.33 and 0.43 [31], according to studies of cured epoxy resin adhesives. The corresponding shear modulus, $G=E/[2(1+\mu)]$, therefore lies in the range of 0.5-1.5 GPa at room temperature. Under the thermal actuation temperature used in this study ($300^{\circ}$F), which exceeds the glass transition temperature of epoxy adhesives, the cured adhesive undergoes pronounced thermal softening and exhibits a substantial reduction in stiffness, by two to three orders of magnitude [32]. Consequently, the Young’s and shear moduli of the adhesive during thermal actuation are on the order of MPa, rather than GPa.
Incorporating such a thin adhesive layer would effectively lead to a trilayer composite with slightly increased effective stretching and bending stiffnesses $k_{s}$ and $k_{b}$ in Equations 18 and 19. However, because the adhesive layer is thin compared to the composite, its primary effect would be a quantitative increase in stiffness rather than a qualitative change in deformation shape.
Importantly, the focus of this work is on the coupling between in-plane stretching and out-of-plane bending induced by thermal strain mismatch in the composite. This coupling mechanism, as well as the resulting morphing behavior, is not sensitive to small variations in effective stiffness caused by the thin adhesive layer. We therefore neglect the adhesive layer in the numerical modeling without loss of generality, while noting that explicitly including it would not alter the main conclusions of the study.
2.3 Numerical Results Validated with Experiments
2.3.1 Model Parameters
The bilayer composite is modeled as a single-layer elastic shell that incorporates thermal strains and dihedral angles as reference strains in the energy, Equation 6. Each region is characterized by its effective Young’s modulus $Y$. For thin plates and shells, it is reasonable to assume plane stress condition, under which the effective Young’s modulus is given by $Y=Y_{0}/(1-\nu^{2})$, where $Y_{0}$ is the original modulus value. At room temperature, for the Shrinky Dink sheet, denoted as layer 1, $Y_{0,1}$ typically ranges from 0.8GPa to 3GPa [33]. For the 3D-printed PLA layer, denoted as layer 2, $Y_{0,2}$ ranges from 2GPa to 4GPa [15]. Without loss of generality, we take the Poisson’s ratio as $\nu=0.33$ for both layers. We set the original modulus $Y_{0,1}=0.8911$ GPa for the Shrinky Dink, and $Y_{0,2}=2.6733$ GPa for the PLA layer, giving effective moduli $Y_{1}=1$ GPa and $Y_{2}=3$ GPa, respectively. At a temperature over $300^{\circ}F$ which is required for thermal actuation, the moduli decrease by two to three orders of magnitude compared to their values at room temperature [30, 46, 38]. So we set $Y_{1}=1$ MPa for the Shrinky Dink, and $Y_{2}=3$ MPa for the PLA layer. Given the scaling of stiffness and energy are both linear in Young’s modulus, the unit from GPa to MPa does not affect the equilibrium solution. The effective stretching stiffness for a bilayer composite is [49]
@@ 2.3.1 Model Parameters
$D/b=D_{\rm eff}/b=\sum_{i=1}^{2}Y_{i}\biggl[\frac{h_{i}^{3}}{12}+h_{i}(y_{i}-\bar{y})^{2}\biggr],$ (19)
where the location of the bending neutral axis is $\bar{y}=(\sum_{i=1}^{2}Y_{i}A_{i}y_{i})/(\sum_{i=1}^{2}Y_{i}A_{i})$ [49]. We set the bottom of Shrinky Dink (layer 1) at $y=0$, so we have $y_{1}=h_{1}/2$, $y_{2}=h_{1}+h_{2}/2$, $A_{1}=bh_{1}$, and $A_{2}=bh_{2}$.
@@
Recall that layer 1 in Equation 19 denotes the Shrinky Dink sheet, and layer 2 denotes the PLA layer. In our experiments, we used thicknesses of $h_{1}=0.3$mm for the Shrinky Dink sheet, and $h_{2}=0.7$ $m$, $0.6$ mm, and 1.01.0mm for the PLA layer in patterns A, B, and C in Figure 2, respectively.
From our experimental observation, bending is activated more easily than in-plane stretching during the thermal actuation process. To improve numerical stability, we scale the effective stretching stiffness $k_{s}$ up by one order of magnitude. This scaling is equivalent to reducing the effective bending stiffness $k_{b}$ by the same factor, since the elastic energy density in Equation 6 depends linearly on both stiffnesses. We have verified numerically that both approaches lead to identical equilibrium configurations.
For each layer, Equation 5 gives the ratio between stretching stiffness $k_{s}$ and bending stiffness $k_{b}$ as
$\frac{k_{s}}{k_{b}}=9\frac{l^{2}_{0}}{h^{2}}.$ (20)
For a thin shell with $l_{0}/h<1$, the bending stiffness is orders of magnitude smaller than the stretching stiffness. As a result, in classical thin shell problems with free bending, stretching energy usually dominates bending energy due to the relatively large ratio $k_{s}/k_{b}$. However, in the present thermally actuated system, the deformation is governed by an imposed in-plane strain mismatch. As the structure bends to accommodate the prescribed thermal strain, the stretching term contributing to the stretching energy, $\varepsilon-\varepsilon^{\rm th}$, remains small, and the dominant contribution arises from bending. This behavior is characteristic of curvature-dominated responses.
We choose to scale $k_{s}$ rather than $k_{b}$ because stretching contributes less to the total strain energy than bending for the geometries and boundary conditions considered here. We have verified that scaling $k_{s}$ upward or scaling $k_{b}$ downward by the same factor leads to identical solutions. However, when $k_{s}$ is scaled by smaller factors, such as 2 or 5, some regions may favor stretching over bending, leading to different equilibrium configurations. Scaling $k_{s}$ by one order of magnitude consistently generates stable solutions.
@@
We solve for equilibrium at each time step using a Newton-Raphson iteration, which corresponds to locally minimizing the total potential energy under the prescribed boundary conditions. As is common for solvers of nonlinear mechanical systems, the Newton method converges to a nearby stable equilibrium configuration rather than guaranteeing convergence to a global minimum. Thermal actuation is applied in incremental load steps with an adaptive step size chosen to satisfy the convergence criterion $|\Delta\mathbb{X}|/|\mathbb{X}|<10^{-5}$. Thermal strains are prescribed according to Equation 16, see Section 2.2.2 for justification. Because the Shrinky Dink layer contracts on heating, the induced thermal strains are negative in sign, $\varepsilon^{\rm th}_{\rm pre}<0$. To account for practical oven heating and to improve numerical stability, we introduce a small, spatially uniform thermal flux load during the heating process. This load gradually decays to zero by the end of actuation, consistent with the sample being removed from the oven. The small load promotes smooth convergence from the initial flat configuration to a physically meaningful deformation branch, especially in cases where the structure may initially bifurcate to bend either upward or downward.
In experiments, bilayer composites are observed to occasionally bend upward or downward depending on small and often imperceptible manufacturing imperfections. Importantly, however, the resulting 3D shapes are very similar. Our numerical simulations reproduce stable 3D configurations for a given composite design and thermal actuation process using the current energy minimization approach. The close quantitative agreement between simulations and experiments, with SSIM values above 0.8, indicates that the computed final shapes approximate the experimental configurations.
While a rigorous proof that the numerical solution corresponds to the global minimum is generally not available for nonlinear shell systems with multiple stable equilibrium states, the good consistency with experimental observations suggests that the calculated configurations represent the experimental responses.
@@
We discretize the composite domain into triangular elements, three meshes in the reference configuration are shown in Fig. 4. We have verified the finite-difference approximations of the gradient and Hessian of the total energy. The numerical implementation was further verified with analytical solutions for uniaxial tension and compression of a simple geometry domain, and the deflection of a cantilever beam under an endpoint load. Complex behaviors, such as bulging and necking also emerge from simple deformation modes [55].
To check the dependence of the results on mesh resolution, we performed a mesh sensitivity analysis using three different meshes for each pattern. For example, for pattern A, the composite was discretized using meshes with 863, 970, and 1080 nodes. The mean structural similarity index (SSIM, defined in Section 2.4) between the deformed configurations obtained using the 863- and 970-node meshes is 0.97, and increases to 0.99 between meshes with 970 and 1080 nodes, indicating convergence of the solution with respect to the mesh discretization. Similar convergence was observed in nodal displacement fields. Based on these results, the meshes shown in Figure 4 A-C(i) are used in simulations. Given these verifications, all subsequent simulations are obtained with the same numerical implementations.
2.3.2 Validation of DES Results with Experimental Scans
@@
Figure 5 presents the experimental scans, numerical simulations, contours of dihedral angle $|\theta|$, and axial strains $\varepsilon$ for a circular composite with a six-arm PLA pattern, Figure 2A. The term “stimulus” refers to the monotonically increasing stimulus, which is heat here, during the process. From top to bottom, the simulations are in equilibrium with prescribed thermal strains $\varepsilon^{\rm th}_{\rm pre}=-0.20,-0.30$ and 0.77-0.77, respectively. As noted above, these thermal strains are nonuniform and vary with each material point’s distance to the PLA pattern. Larger thermally induced shrinkage results in more significant deformations and axial strains. A greater mismatch in axial strains leads to larger dihedral angles. To accommodate these deformations, the structure morphs from a planar configuration into a 3D shape. Deformation localizes heavily between the arms of the 3D-printed PLA pattern, where the Shrinky Dink layer undergoes the largest shrinkage, indicated by A-C(iii) and A-C(iv) in Figure 5. Overall, excellent agreement is observed between the scanned experimental shapes (first column) and the numerical simulations at each stage of the thermal actuation.
Figure 5: Representative snapshots, starting with the pattern of Figure 2A during the thermally actuated manufacturing process. From the top to the bottom rows, the duration of the thermal stimulus increases. In each column, (i) shows the experimentally scanned shape, (ii) shows the numerical simulation, (iii) shows the distribution of the absolute value of the dihedral angle $|\theta|$, and (iv) shows the distribution of the axial strain $\varepsilon$.
@@
Figure 6 shows the results for the rectangular composite with a two-star PLA pattern in Figure 2B, subjected to thermal strains $\varepsilon^{\rm th}_{\rm pre}=-0.08,-0.15$ and 0.30-0.30, from the top to the bottom, respectively. As the prescribed thermal axial strain increases over time, the structure develops larger axial strains $\varepsilon$, strains mismatch, dihedral angles $|\theta|$, and overall curvature. Deformation localizes near the four corners of the Shrinky Dink layer. The slight asymmetry in the deformed shapes arises from the mesh used in the numerical simulation and minor anisotropies in the bilayer composite due to physical manufacturing defects. Nevertheless, the numerical simulations closely approximate the scanned experimental results.
@@
Similarly, Figure 7 illustrates the deformed configurations at thermally actuated strains $\varepsilon^{\rm th}_{\rm pre}=-0.10$, 0.30-0.30 and 0.50-0.50 for the cross pattern in Figure 2C. As the prescribed thermal strain increases, deformation concentrates increasingly between pattern’s arms. Because the Shrinky Dink layer is square rather than circular, these concentration zones form four sharp “petals” instead of the rounded bowl seen in Figure 5.
@@
These three examples demonstrate excellent agreement between scanned experimental 3D shapes and the numerical simulations morphed from 2D planar composites across various domain geometries, patterns, and levels of thermal actuation. More complex structures can be achieved by assembling multiple shapes and using multiscale modeling. The substantial computational demand of multiscale modeling across multiple time and length scales can be mitigated through iterated learning of operators for out-of-distribution data [55]. Potential applications of the three examples include deployable flower petals, bowls, frame-based pillows, canoes, and tent-type structures, each achieved by selecting an appropriate reference length scale.
The current model predicts deformation morphology under the assumption of ideal bonding and purely elastic behavior. However, during experiments we observed localized failure phenomena, including interface debonding near the boundary between the kirigami PLA layer and the Shrinky Dink layer, as well as strain localization at sharp corners. Because the thermal actuation temperature exceeds $300^{\circ}$F (shown by Figure 3E), which is higher than the glass transition temperatures of both the Shrinky Dink substrate and the PLA layer, the dominant deformation mechanisms associated with local failures are likely creep deformation, stress relaxation, and viscoelastic effects, rather than brittle fracture or rate-independent yielding. In the absence of significant external loading that would cause tearing, ductile fracture was not observed.
The localized failure phenomena could potentially explain the slight asymmetry in deformation observed experimentally in Figures 5, 6, 7 A(i), B(i) and C(i). At present, the model does not capture such failure mechanisms. Nevertheless, the framework can be extended to incorporate inelastic effects, for example by introducing Perzyna viscoplasticity based on internal variable theory [55], as well as power-law creep models to account for time-dependent creep deformation and stress relaxation [56].
2.3.3 Young’s Modulus and Nominal Stiffness Before and After Thermal Actuation
We further estimate the Young’s modulus of the Shrinky Dink layer before and after thermal actuation. Given the small thickness (approximately $0.3$ mm) of a Shrinky Dink sheet before actuation, we consider a thin cantilever beam with one end fixed with all DOFs constrained, and the other end free. With gravity acting as a distributed load on the beam, the classical relation between the distributed load $q$ and the free-end deflection $\delta_{0}$ is given by
$\delta_{0}=\frac{qL^{4}}{8EI_{0}},\quad I=b_{0}h_{0}^{3}/12,\quad q=\rho gb_{0}h_{0},$ (21)
where $b_{0}$ is the width, $h_{0}$ is the thickness, and the density is $\rho=1.05\rm{g/cm}^{3}$ before thermal actuation. Using this relation, we estimate the Young’s modulus as $E=2.60$ GPa with a standard deviation of $0.32$ GPa before thermal actuation.
After thermal actuation, the Shrinky Dink sheet becomes thicker, with a thickness of approximately $1.8$ mm, and gravity no longer generates a noticeable deformation. In this case, we apply a point load at the free end of the cantilever beam, and measure the applied force using a Rokubi six-axis force sensor (BOTA BFT-ROKA-SER-M8) [10]. Using the classical relation
$\delta_{a}=\frac{PL^{3}}{3EI_{a}},\quad I_{a}=b_{a}h_{a}^{3}/12$ (22)
where the subscript $(\cdot)_{a}$ denotes values after thermal actuation, we estimate the Young’s modulus as $E=2.65$ GPa with a standard deviation of $1.58$ GPa after thermal actuation. The relatively large standard deviation is due to the high nominal stiffness of the thickened Shrinky Dink layer after thermal actuation, which makes accurate deformation measurements more challenging. The reported results are averaged over five samples in each case.
The nominal stiffness of a complex composite is an important structural performance metric, particularly in applications such as aerospace engineering. The nominal stiffness $K_{\rm nom}$ of a structure is defined as the ratio between the applied generalized load $P_{\rm nom}$ and the measured generalized displacement $\delta_{\rm nom}$. For a cantilever beam with a point load $P$ applied at the free end, leading to a displacement $\delta_{a}$, we set $P_{\rm nom,a}=P$, $\delta_{\rm nom,a}=\delta_{a}$, and the nominal stiffness is given by
$K_{\rm nom,a}=P_{\rm nom,a}/\delta_{\rm nom,a}=3EI_{a}/L^{3},$ (23)
from Equation (22). For a cantilever beam subjected to a distributed load $q$ with free-end displacement $\delta_{0}$, we set $P_{\rm nom,0}=qL$ and $\delta_{\rm nom}=\delta_{0}$, which gives
$K_{\rm nom,0}=P_{\rm nom,0}/\delta_{\rm nom,0}=8EI_{0}/L^{3},$ (24)
from Equation (21). Here, the moment of inertia $I$, which changes significantly due to geometric thickening after thermal actuation, dominates the nominal stiffness. Using $I_{0}~10^{-14}m^{4}$ and $I_{a}~10^{-12}m^{4}$, along with $L_{0}=0.16$ m and $L_{a}~0.073$ m, we obtain $K_{\rm nom,a}=142\pm 52$ N/m, and $K_{\rm nom,0}=0.2357\pm 0.0016$ N/m. It is worth noting that the estimated Young’s modulus values for the Shrinky Dink material before and after thermal actuation remain close at approximately 2.6 GPa, whereas the corresponding nominal stiffness values differ by more than a factor of 600.
These results indicate that, although the Young’s modulus of the Shrinky Dink material remains essentially unchanged before and after thermal actuation, the rigidity of the cooled, permanently deformed structure increases dramatically due to geometric stiffening, resulting in a nominal stiffness more than two orders of magnitude larger than that of the undeformed configuration.
2.4 Structural Similarity Index Measure